refer to the stock options on Apple

O No No Negative value shouid be
media%2F3c3%2F3c35d44b-8617-4b8d-8341-65
Answer #1

a-1. Yes, the strike price is lesser than the stock price which means exercising call would be cheaper than buying in market. However the call premium is $2.62.

a-2. There is net loss of $(102-100)-$2.62 x 100 shares= $62

a-3. The rate of return is -0.62/100= -0.62%

b-1. Yes.

b-2. Net profit= $(102-95)-$6.35= $0.65

b-3. Rate of return= 0.65/95= 0.68%

c-1. No, exercising put at a stock price of 100 will lead to a 0 profit situation.

c-2.There is a premium loss of $1.55 per share. Rate of return is 1.55/100= -1.55%